OpenStudy (anonymous):

The weight of an object on the moon varies directly as its weight on Earth. With all of his gear on, Neil Armstrong weighed 360 pounds on Earth. When he became the first person to step on the moon on July 20, 1969, he weighed 60 pounds. Tara weighs 108 pounds on Earth. What would she weigh on the moon? help P L E A S E

OpenStudy (anonymous):

18 pounds

OpenStudy (anonymous):

W(Earth) = k x W(Moon) Plugging Neil Armstrong's weight values on earth and moon we can obtain the value of k/

OpenStudy (amistre64):

e = k.m 360 = k.60 k = 360/60 = 36/6 = 6 108 = 6.m 108/6 = m = 18 ; him was right :)

OpenStudy (anonymous):

thanks (: (: (: (: ya'll are A M A Z I N G! <3

OpenStudy (amistre64):

well, him is amazing maybe; im just average lol

OpenStudy (anonymous):

so many compliments///its lucky its dark..i havent blushed so much since madam pomfrey told me she liked my new earmuffs...:)

OpenStudy (anonymous):

amistre64 your amazing toooo!!! and him1618 lol your welcome :D

OpenStudy (anonymous):

u got where i picked that from..m a big dumbledore fan.thnx a lot

OpenStudy (anonymous):

oh cool!!